what is the distance between the points (3,3) and (2,1)?
Answer by checkley75(3666)

the hypotenuse of the triangle formed by these two points is:
h^2=(3-2)^2+(3-1)^2
h^2=1^2+2^2
h^2=1+4
h^2=5
x=sqrt5
h=2.236 is the distance between these two points.